# Scaffold-Aware Generative Augmentation and Reranking for Enhanced Virtual Screening

### v1

@@ Abstract
Ligand-based virtual screening (VS) is an essential step in drug discovery that evaluates large chemical libraries to identify compounds that potentially bind to a therapeutic target. However, VS faces three major challenges: class imbalance due to the low active rate, structural imbalance among active molecules where certain scaffolds dominate, and the need to identify structurally diverse active compounds for novel drug development. We introduce ScaffAug, a scaffold-aware VS framework that addresses these challenges through three modules. The augmentation module first generates synthetic data conditioned on scaffolds of actual hits using generative AI, specifically a graph diffusion model. This helps mitigate the class imbalance and furthermore the structural imbalance, due to our proposed scaffold-aware sampling algorithm, designed to produce more samples for active molecules with underrepresented scaffolds. A model-agnostic self-training module is then used to safely integrate the generated synthetic data from our augmentation module with the original labeled data. Lastly, we introduce a reranking module that improves VS by enhancing scaffold diversity in the top recommended set of molecules, while still maintaining and even enhancing the overall general performance of identifying novel, active compounds. We conduct comprehensive computational experiments across five target classes, comparing ScaffAug against existing baseline methods by reporting the performance of multiple evaluation metrics and performing ablation studies on ScaffAug. Overall, this work introduces novel perspectives on effectively enhancing VS by leveraging generative augmentations, reranking, and general scaffold-awareness.
@@ 1 Introduction
Virtual screening (VS) supports traditional drug discovery through computational identification of promising compounds for experimental validation, reducing time, cost, and labor (Leelananda & Lindert, 2016). A fundamental task in VS involves predicting the biological activity of chemical compounds against specific therapeutic targets based on their molecular structures. Recent advances in deep learning methods, particularly graph neural networks (GNNs), have significantly improved the effectiveness of these predictive tasks (Golkov et al., 2020). However, three key challenges remain. The first challenge is class imbalance, which occurs because large libraries of molecules are tested in high-throughput screening (HTS), but only a small percentage are actual actives (Zhu et al., 2013). This means limited known, active molecules are available for reference, making model training difficult and biasing predictions toward the far more numerous inactive molecules. The second challenge is structural imbalance among known active molecules in the training data. This imbalance among known active molecules is commonly observed in VS datasets like WelQrate (Liu et al., 2024b), where some groups of similar active molecules form dominant clusters. Meanwhile, many other active molecules possess structures significantly different from those in the clusters, making them underrepresented in the data. As a result, models become biased toward dominant structural clusters, potentially overlooking novel candidates with distinct molecular structures. The third challenge arises from the requirement in actual drug discovery applications to discover novel molecules distinct from existing patented structures. To increase the probability of finding novel active compounds across different regions of chemical space, medicinal chemists maximize scaffold diversity when constructing compound libraries for HTS (Langdon et al., 2011) and creating virtual screening libraries (Krier et al., 2006). This involves clustering molecules based on their core structures (scaffolds) and selecting representative compounds from each cluster, with protocols like DEKOIS 2.0 (Bauer et al., 2013), which selects the most potent compound from distinct structural clusters. Deep learning models often struggle to identify novel active molecules because they tend to overfit and prioritize compounds with structural features similar to active molecules in the training data, rather than recognizing diverse structural patterns that may still produce biological activity.
@@
We propose to leverage synthetic data generation (Lu et al., 2023) through generative AI, specifically a graph diffusion model (Vignac et al., 2022) for small molecules to address both class and structural imbalance challenges. This method creates new samples conditioned on known active molecules’ scaffolds to generate likely active compounds, effectively balancing the representation of active versus inactive molecules in the training data. By directing the generation process toward underrepresented molecular scaffolds via our proposed scaffold-aware sampling, we produce sufficient samples that ensure adequate representation of minor structural families, enabling models to learn more comprehensive structure-activity relationships beyond dominant structural clusters. We note that this approach overcomes the inherent challenges of directly generating molecules and instead integrates the benefits of molecular generation into our enhanced virtual screening framework. To tackle the third challenge of discovering novel molecules, we design a reranking algorithm that injects scaffold diversity consideration into deep learning model’s top prediction set output. This approach helps identify promising candidates with structural novelty that models’ predicted activity scores might otherwise overlook.
@@
In the reranking module, we apply Maximal Marginal Relevance (MMR) (Carbonell & Goldstein, 1998) to encourage scaffold diversity in the model’s top predicted set. This approach balances the model’s predicted scores with a diversity score, reranking the list to achieve higher scaffold diversity while maintaining screening quality.
@@ 2.2 Related Work
Graph-level augmentation improves training on imbalanced data by adding auxiliary information (Ma et al., 2025). Examples include G-mixup, which interpolates graph generators (graphons) across classes (Han et al., 2022), and G2GNN, which builds a graph-of-graphs to enrich minority class context (Wang et al., 2022). FLAG (Kong et al., 2022) perturbs node features adversarially, while GREA (Liu et al., 2022) augments subgraphs. However, such general-purpose methods ignore chemical valency rules, often producing invalid molecules that hinder training. G2GNN is further limited by scalability, as it requires pairwise similarity across all samples. By contrast, DCT (Liu et al., 2023a), a chemistry-specific diffusion model, generates valid task-specific examples, though its benefits for virtual screening remain uncertain given the extreme class imbalance discussed in Sec. 1.
@@
AI-driven molecule design spans variational autoencoders (Blaschke et al., 2018), BERT-style models (Wang et al., 2019), GANs (De Cao & Kipf, 2018), and graph neural networks (Li et al., 2018). Recent work has adapted GPT-like language models for molecular generation (Bagal et al., 2021; ). Diffusion models, first developed for image synthesis (Ho et al., 2020), are now widely used across domains, including molecular graphs, where they address permutation invariance and discrete structure (Niu et al., 2020; Vignac et al., 2022). Conditional variants guide generation by imposing structural or property constraints (Huang et al., 2023; Liu et al., 2024a). In our work, we employ DiGress (Vignac et al., 2022) to generate molecules conditioned on scaffolds of known actives, thus steering synthesis toward chemically meaningful regions. Integrated into our augmentation framework, this conditional approach produces structurally diverse molecules and helps alleviate both class and structural imbalance in VS datasets.
@@ 3.1 Augmentation Module: Mitigating Class and Scaffold Imbalances
The augmentation module, as shown in Figure 2(a), generates synthetic data to address the class imbalance by increasing the proportion of active molecules. To ensure the synthetic data remains within the active molecule distribution, we preserve the scaffolds of active compounds and generate new molecules with a graph diffusion model, a process we call scaffold extension. Before this step, we use a scaffold-aware sampling algorithm to balance the dominant and underrepresented active scaffolds when building the scaffold library, tackling the structural imbalance. known, active molecules in the train set by prioritizing underrepresented scaffolds for scaffold extension.
Scaffold-aware Sampling (SAS) Algorithm. We begin by extracting scaffold SMILES strings from active molecules in the training set. These scaffolds are then converted to 1024-bit Extended-Connectivity Fingerprints (ECFP)(Rogers & Hahn, 2010) to enable structural similarity measurement. We apply K-Means(MacQueen, 1967) clustering to assign scaffold labels, with the optimal number of clusters determined by Silhouette scores(Rousseeuw, 1987). Our sampling procedure calculates weights inversely proportional to cluster frequencies, giving higher priority to less common structural patterns. After normalizing these weights into sampling probabilities, we select scaffolds with replacement using these probabilities, where equals 10% of the training data size. We denote the resulting scaffold graph library as , which serves as the foundation for the scaffold extension process. Details of the algorithm are provided in Alg. 1.
@@ 3.3 Reranking Module: Preserving Scaffold Diversity in the Top-ranked Set
Virtual screening is typically handled as a ranking problem because pharmaceutical companies only purchase or synthesize the top-ranked molecules, while those with lower prediction scores receive less attention. To tackle the challenge of identifying novel active compounds, we design a Maximal Marginal Relevance (MMR) (Carbonell & Goldstein, 1998) reranking algorithm to enhance the structural diversity within the top-ranked set initially output by the VS model .
@@ 4.1 Experimental Setup
Datasets. For this work, we utilize a subset of the comprehensive WelQrate(Liu et al., 2024b) dataset is a high-quality dataset of 9 bioassays covering five therapeutic target classes for small-molecule drug discovery benchmarking. Its quality stems from a rigorous hierarchical curation process that includes primary screens followed by confirmatory and counter screens, along with extensive filtering(e.g., PAINS filtering, druglikeness assessment) and expert verification. The datasets realistically represent drug discovery campaigns with large compound numbers (66K-300K) and low active percentages ( 1%). Each WelQrate dataset possesses ten split schemes for robust evaluation. Five random splits adopt a nested cross-validation strategy to allow every molecule in the dataset to be evaluated in the test set. Five scaffold splits with different random seeds separate molecules with different core structures to test models’ ability to predict activity across structural variations. More detailed discussion on the random vs scaffold splits is in Appendix A.3 We fine-tuned different methods on each dataset’s split and reported the aggregated result for random and scaffold splits, respectively. For this work, we select five representative datasets that cover all therapeutic target types in the WelQrate dataset for evaluation; dataset details are shown in Appendix A.1.
@@
BEDROC ranges from 0 to 1, where a score closer to 1 indicates better performance in recognizing active compounds early in the list (Pearlman & Charifson, 2001).
@@
DCG100 aims to penalize a true active molecule appearing lower in the selection set by logarithmically reducing the relevance value proportional to the predicted rank of the compound within the top 100 predictions (Järvelin & Kekäläinen, 2017).
We present the formal definitions for each of the above along with more detailed justification for their use in benchmarking in Appendix A.4.
Furthermore, we include a novel scaffold diversity metric SD100 designed to quantify the dissimilarity between molecular scaffolds through pairwise Tanimoto similarity of ECFP fingerprints:
@@
Baselines and implementation. We implement various graph neural networks (GNNs) as backbone models to demonstrate the generalizability of ScaffAug, including GCN(Kipf & Welling, 2017), GAT(Veličković et al., 2018), and DeepGIN(Xu et al., 2019). We also compare three recent graph augmentation methods: FLAG(Kong et al., 2022), GREA(Liu et al., 2022), and DCT(Liu et al., 2023a), which show potent performance on molecule property prediction. To ensure fair comparison on the severely class-imbalanced WelQrate dataset, we apply an Imbalanced Sampler for ScaffAug and all baseline methods. Therefore, we mark the baseline with os, indicating that oversampling of the minor class is used. Each baseline is carefully fine-tuned based on the hyperparameter pool given by its original paper and code with the Optuna package(Akiba et al., 2019). During training, we use BEDROC as the validation metric due to its stability, which helps produce test results with lower variance. The epoch with the highest BEDROC score on the validation set is selected for the final testing evaluation. We conduct three testing evaluations with different random seeds for each split scheme and aggregate results for the random and scaffold splitting strategies separately.
@@ 4.2 Experimental Results
Reranking Analysis. After evaluating the trained virtual screening model on the test set, we derive prediction scores for all test molecules. We only keep the positive scores and sort them, forming the candidate set with a maximum size of 500 molecules. The next step applies MMR reranking to to obtain the reranked set . To assess the quality of this reranking, the enrichment factor EF100 and the scaffold diversity SD100 are calculated on the top 100 subset of the original and reranked set . Table 3 presents the reranking results of ScaffAug across all five WelQrate datasets, showing SD100 and EF100 for reranked sets with different values for MMR score calculation. With smaller values, we gain greater scaffold diversity in the set. Conversely, when , MMR selects compounds based on the original prediction scores without considering structural diversity. Notably, when scaffold diversity is overemphasized (), EF100 typically decreases, indicating that careful tuning of is required for each dataset. For the AID1798 and AID2689 scaffold splits, where the model’s ability to generalize across structural variations is more severely tested, EF100 increases when more scaffold diversity is incorporated into . Specifically, for the AID1798 scaffold split, EF100 increases by 19.8% while SD100 increases by 2%. In other cases, reranking can maintain EF100 while increasing some scaffold diversity. This motivates us to practice similar reranking methods focused on structural diversity in the actual virtual screening scenarios.
@@
Ablation Study of Augmentation Module. The augmentation module explores the chemical space around active molecules in the training set, providing additional knowledge for the limited active samples to strengthen model performance. We carefully examine the usefulness of our scaffold-aware sampling (SAS) algorithm on the structural imbalance among active molecules in the training set. As shown in Fig. 4, the active molecules in the training set (AID1798) exhibit a dominant cluster in the central region of the UMAP projection, with over half of the active molecules concentrated in this area, while the remaining molecules have limited presence in peripheral chemical regions. We compare SAS with the uniform actives (UA) sampling algorithm, which selects each active scaffold uniformly with equal probability in the augmentation module. With UA sampling, the generated molecules maintain a similar distribution pattern to the original training data, with most molecules clustering in the central region and leaving peripheral chemical space under-explored, as evidenced by molecular counts reaching over 250 in the central region while maintaining near-zero representation in peripheral areas. This concentration bias makes the identification of novel active molecules in peripheral areas challenging, as the model receives insufficient training examples from these regions. In contrast, SAS addresses this structural imbalance by assigning higher sampling weights to underrepresented scaffold structures when constructing the scaffold library for molecule generation. The resulting SAS-generated molecules demonstrate improved spatial distribution across the UMAP projection, with more uniform coverage of peripheral chemical regions that were undersampled by UA. This enhanced coverage results in a G-DSA dataset that better represents the full chemical diversity around active compounds in the training set, providing more balanced training examples across different structural motifs and chemical space regions.
Figure 5 compares the percentage change of scaffold-aware sampling (SAS) and uniform actives (UA) variants relative to the oversampling (os) baseline. ScaffAug with SAS consistently outperforms both the oversampling baseline and the UA variant across all evaluation metrics, with this improvement pattern observed across different model architectures.
@@
This consistency indicates the effectiveness and robustness of the scaffold-aware generative augmentation framework for virtual screening tasks. The SAS approach achieves particularly substantial improvements in EF100 and DCG100 metrics, with performance gains reaching approximately 45% and 38% respectively in optimal cases. These enrichment-based metrics are critical indicators of early retrieval performance in drug discovery workflows. The UA variant shows markedly smaller improvements for these same metrics and, more significantly, performs worse than the oversampling baseline for EF100 (declining over 10%) and DCG100 metrics when applied to certain model architectures. This performance degradation occurs because uniform sampling of active compounds worsen structural imbalance by generating additional molecules from already over-represented scaffold clusters. The resulting bias toward dominant structural patterns impairs the model’s ability to generalize to structurally distinct molecules during evaluation. These results underscore the importance of scaffold-aware sampling strategies that actively increase scaffold diversity among active compounds in the training data, rather than simply augmenting the total number of active samples.

### v2

@@ Abstract
Ligand-based virtual screening (VS) is an essential step in drug discovery that evaluates large chemical libraries to identify compounds that potentially bind to a therapeutic target. However, VS faces three major challenges: class imbalance due to the low active rate, structural imbalance among active molecules where certain scaffolds dominate, and the need to identify structurally diverse active compounds for novel drug development. We introduce ScaffAug, a scaffold-aware VS framework that addresses these challenges through three modules. The augmentation module first generates synthetic data conditioned on scaffolds of actual hits using generative models, specifically a graph diffusion model. This helps mitigate the class imbalance and furthermore the structural imbalance, due to our proposed scaffold-aware sampling algorithm, designed to produce more samples for active molecules with underrepresented scaffolds. A model-agnostic self-training module is then used to safely integrate the generated synthetic data from our augmentation module with the original labeled data. Lastly, we introduce a reranking module that improves VS by enhancing scaffold diversity in the top recommended set of molecules, while still maintaining and even enhancing the overall general performance of identifying novel, active compounds. We conduct comprehensive computational experiments across five target classes, comparing ScaffAug against existing baseline methods by reporting the performance of multiple evaluation metrics and performing ablation studies on ScaffAug. Overall, this work introduces novel perspectives on effectively enhancing VS by leveraging generative augmentations, reranking, and general scaffold-awareness.
@@ 1 Introduction
Virtual screening (VS) supports traditional drug discovery through computational identification of promising compounds for experimental validation, reducing time, cost, and labor (Leelananda and Lindert, 2016). A fundamental task in VS involves predicting the biological activity of chemical compounds against specific therapeutic targets based on their molecular structures. Recent advances in deep learning methods, particularly graph neural networks (GNNs), have significantly improved the effectiveness of these predictive tasks (Golkov et al., 2020). However, three key challenges remain. The first challenge is class imbalance, which occurs because large libraries of molecules are tested in high-throughput screening (HTS), but only a small percentage are actual actives (Zhu et al., 2013). This means limited known, active molecules are available for reference, making model training difficult and biasing predictions toward the far more numerous inactive molecules. The second challenge is structural imbalance among known active molecules in the training data. This imbalance among known active molecules is commonly observed in VS datasets like WelQrate (Liu et al., 2024b), where some groups of similar active molecules form dominant clusters. Meanwhile, many other active molecules possess structures significantly different from those in the clusters, making them underrepresented in the data. As a result, models become biased toward dominant structural clusters, potentially overlooking novel candidates with distinct molecular structures. The third challenge arises from the requirement in actual drug discovery applications to discover novel molecules distinct from existing patented structures. To increase the probability of finding novel active compounds across different regions of chemical space, medicinal chemists maximize scaffold diversity when constructing compound libraries for HTS (Langdon et al., 2011) and creating virtual screening libraries (Krier et al., 2006). This involves clustering molecules based on their core structures (scaffolds) and selecting representative compounds from each cluster, with protocols like DEKOIS 2.0 (Bauer et al., 2013), which selects the most potent compound from distinct structural clusters. Deep learning models often struggle to identify novel active molecules because they tend to overfit and prioritize compounds with structural features similar to active molecules in the training data, rather than recognizing diverse structural patterns that may still produce biological activity.
@@
We propose to leverage synthetic data generation (Lu et al., 2023) through generative models, specifically a graph diffusion model (Vignac et al., 2022) for small molecules to address both class and structural imbalance challenges. This method creates new samples conditioned on known active molecules’ scaffolds to generate likely active compounds, effectively balancing the representation of active versus inactive molecules in the training data. By directing the generation process toward underrepresented molecular scaffolds via our proposed scaffold-aware sampling, we produce sufficient samples that ensure adequate representation of minor structural families, enabling models to learn more comprehensive structure-activity relationships beyond dominant structural clusters. We note that this approach overcomes the inherent challenges of directly generating molecules and instead integrates the benefits of molecular generation into our enhanced virtual screening framework. To tackle the third challenge of discovering novel molecules, we design a reranking algorithm that injects scaffold diversity consideration into deep learning model’s top prediction set output. This approach helps identify promising candidates with structural novelty that models’ predicted activity scores might otherwise overlook.
@@
In the reranking module, we apply Maximal Marginal Relevance (MMR) (Carbonell and Goldstein, 1998) to encourage scaffold diversity in the model’s top predicted set. This approach balances the model’s predicted scores with a diversity score, reranking the list to achieve higher scaffold diversity while maintaining screening quality.
@@ 2.2 Related Work
Graph-level augmentation improves training on imbalanced data by adding auxiliary information (Ma et al., 2025). Examples include G-mixup, which interpolates graph generators (graphons) across classes (Han et al., 2022), and G2GNN, which builds a graph-of-graphs to enrich minority class context (Wang et al., 2022). FLAG (Kong et al., 2022) perturbs node features adversarially, while GREA (Liu et al., 2022) augments subgraphs. However, such general-purpose methods ignore chemical valency rules, often producing invalid molecules that hinder training. G2GNN is further limited by scalability, as it requires pairwise similarity across all samples. By contrast, DCT (Liu et al., 2023a), a chemistry-specific diffusion model, generates valid task-specific examples, though its benefits for virtual screening remain uncertain given the extreme class imbalance discussed in Sec. 1. InstructMol (Cao et al., 2025) introduces a novel semi-supervised framework that improves molecular property prediction by using a target model to generate pseudo-labels for unlabeled molecules and an instructor model to evaluate their reliability.
@@
AI-driven molecule design spans variational autoencoders (Blaschke et al., 2018), BERT-style models (Wang et al., 2019), GANs (De Cao and Kipf, 2018), and graph neural networks (Li et al., 2018). Recent work has adapted GPT-like language models for molecular generation (Bagal et al., 2021; ). Diffusion models, first developed for image synthesis (Ho et al., 2020), are now widely used across domains, including molecular graphs, where they address permutation invariance and discrete structure (Niu et al., 2020; Vignac et al., 2022). Conditional variants guide generation by imposing structural or property constraints (Huang et al., 2023; Liu et al., 2024a). In our work, we employ DiGress (Vignac et al., 2022) to generate molecules conditioned on scaffolds of known actives, thus steering synthesis toward chemically meaningful regions. Integrated into our augmentation framework, this conditional approach produces structurally diverse molecules and helps alleviate both class and structural imbalance in VS datasets.
@@ 3.1 Augmentation Module: Mitigating Class and Scaffold Imbalances
The augmentation module, as shown in Figure 2(a), generates synthetic data to address the class imbalance by increasing the proportion of active molecules. To ensure the synthetic data remains within the active molecule distribution, we preserve the scaffolds of active compounds and generate new molecules with a graph diffusion model, a process we call scaffold extension. Before this step, we use a scaffold-aware sampling(SAS) algorithm to balance the dominant and underrepresented active scaffolds when building the scaffold library, tackling the structural imbalance shown in Fig. 5. We prioritize underrepresented scaffolds from the known, active molecules for the scaffold extension.
Scaffold-aware Sampling (SAS) Algorithm. We begin by extracting scaffold SMILES strings from active molecules in the training set. These scaffolds are then converted to 1024-bit Extended-Connectivity Fingerprints (ECFP)(Rogers and Hahn, 2010) to enable structural similarity measurement. We apply K-Means(MacQueen, 1967) clustering to assign scaffold labels, with the optimal number of clusters determined by Silhouette scores(Rousseeuw, 1987). Our sampling procedure calculates weights inversely proportional to cluster frequencies, giving higher priority to less common structural patterns. After normalizing these weights into sampling probabilities, we select scaffolds with replacement using these probabilities, where equals 10% of the training data size. We denote the resulting scaffold graph library as , which serves as the foundation for the scaffold extension process. Details of the algorithm are provided in Alg. 1.
@@ 3.3 Reranking Module: Preserving Scaffold Diversity in the Top-ranked Set
Virtual screening is typically handled as a ranking problem because pharmaceutical companies only purchase or synthesize the top-ranked molecules, while those with lower prediction scores receive less attention. To tackle the challenge of identifying novel active compounds, we design a Maximal Marginal Relevance (MMR) (Carbonell and Goldstein, 1998) reranking algorithm to enhance the structural diversity within the top-ranked set initially output by the VS model .
@@ 4.1 Experimental Setup
Datasets. For this work, we utilize the comprehensive WelQrate (Liu et al., 2024b) dataset, which is a high-quality collection of nine bioassays covering five therapeutic target classes for small-molecule drug discovery benchmarking. Its quality stems from a rigorous hierarchical curation process that includes primary screens followed by confirmatory and counter screens, along with extensive filtering(e.g., PAINS filtering, druglikeness assessment) and expert verification. The datasets realistically represent drug discovery campaigns with large compound numbers (66K-300K) and low active percentages ( 1%). Each WelQrate dataset possesses ten split schemes for robust evaluation. Five random splits adopt a nested cross-validation strategy to allow every molecule in the dataset to be evaluated in the test set. Five scaffold splits with different random seeds separate molecules with different core structures to test models’ ability to predict activity across structural variations. More detailed discussion on the random vs scaffold splits is in Appendix A.3 We fine-tuned different methods on each dataset’s split and reported the aggregated result for random and scaffold splits, respectively. For this work, we select five representative datasets, shown in Appendix A.1, that cover all therapeutic target types in the WelQrate dataset for evaluation.
@@
BEDROC ranges from 0 to 1, where a score closer to 1 indicates better performance in recognizing active compounds early in the list (Pearlman and Charifson, 2001).
@@
DCG100 aims to penalize a true active molecule appearing lower in the selection set by logarithmically reducing the relevance value proportional to the predicted rank of the compound within the top 100 predictions (Järvelin and Kekäläinen, 2017).
We present the formal definitions for each of the above along with more detailed justification for their use in benchmarking in Appendix A.4. Furthermore, we include a novel scaffold diversity metric SD100 designed to quantify the dissimilarity between molecular scaffolds through pairwise Tanimoto similarity of ECFP fingerprints:
@@
Baselines and implementation. We implement various graph neural networks (GNNs) as backbone models to demonstrate the generalizability of ScaffAug, including GCN(Kipf and Welling, 2017), GAT(Veličković et al., 2018), and DeepGIN(Xu et al., 2019). We also compare three recent graph augmentation methods: FLAG(Kong et al., 2022), GREA(Liu et al., 2022), and DCT(Liu et al., 2023a), which show potent performance on molecule property prediction. To ensure fair comparison on the severely class-imbalanced WelQrate dataset, we apply an Imbalanced Sampler for ScaffAug and all baseline methods. Therefore, we mark the baseline with os, indicating that oversampling of the minor class is used. Each baseline is carefully fine-tuned based on the hyperparameter pool given by its original paper and code with the Optuna package(Akiba et al., 2019). During training, we use BEDROC as the validation metric due to its stability, which helps produce test results with lower variance. The epoch with the highest BEDROC score on the validation set is selected for the final testing evaluation. We conduct three testing evaluations with different random seeds for each split scheme and aggregate results for the random and scaffold splitting strategies separately.
@@ 4.2 Experimental Results
Reranking Analysis. After evaluating the trained virtual screening model on the test set, we derive prediction scores for all test molecules. We only keep the positive scores and sort them, forming the candidate set with a maximum size of 500 molecules. The next step applies MMR reranking to to obtain the reranked set . To assess the quality of this reranking, the enrichment factor EF100 and the scaffold diversity SD100 are calculated on the top 100 subset of the original and reranked set . Figure 3 presents the reranking results of ScaffAug across all five WelQrate datasets, showing SD100 and EF100 for reranked sets with different values for MMR score calculation. With smaller values, we gain greater scaffold diversity in the set. Conversely, when , MMR selects compounds based on the original prediction scores without considering structural diversity. Notably, when scaffold diversity is overemphasized (), EF100 typically decreases, indicating that careful tuning of is required for each dataset. For the AID1798 and AID2689 scaffold splits, where the model’s ability to generalize across structural variations is more severely tested, EF100 increases when more scaffold diversity is incorporated into . Specifically, for the AID1798 scaffold split, EF100 increases by 19.8% while SD100 increases by 2%. In other cases, reranking can maintain EF100 while increasing some scaffold diversity. This motivates us to practice similar reranking methods focused on structural diversity in the actual virtual screening scenarios.
Ablation Studies.
Augmentation Module
@@
The augmentation module explores the chemical space around active molecules in the training set, providing additional knowledge for the limited active samples to strengthen model performance. We carefully examine the usefulness of our scaffold-aware sampling (SAS) algorithm on the structural imbalance among active molecules in the training set. As shown in Fig. 5, the active molecules in the training set (AID1798) exhibit a dominant cluster in the central region of the UMAP projection, with over half of the active molecules concentrated in this area, while the remaining molecules have limited presence in peripheral chemical regions. We compare SAS with the uniform actives (UA) sampling algorithm, which selects each active scaffold uniformly with equal probability in the augmentation module. With UA sampling, the generated molecules maintain a similar distribution pattern to the original training data, with most molecules clustering in the central region and leaving peripheral chemical space under-explored, as evidenced by molecular counts reaching over 250 in the central region while maintaining near-zero representation in peripheral areas. This concentration bias makes the identification of novel active molecules in peripheral areas challenging, as the model receives insufficient training examples from these regions. In contrast, SAS addresses this structural imbalance by assigning higher sampling weights to underrepresented scaffold structures when constructing the scaffold library for molecule generation. The resulting SAS-generated molecules demonstrate improved spatial distribution across the UMAP projection, with more uniform coverage of peripheral chemical regions that were undersampled by UA. This enhanced coverage results in a G-DSA dataset that better represents the full chemical diversity around active compounds in the training set, providing more balanced training examples across different structural motifs and chemical space regions.
Self-training Module
@@
Figure 4 compares the percentage change of scaffold-aware sampling (SAS) and uniform actives (UA) variants relative to the oversampling (os) baseline. ScaffAug with SAS consistently outperforms both the oversampling baseline and the UA variant across all evaluation metrics, with this improvement pattern observed across different model architectures. This consistency indicates the effectiveness and robustness of the scaffold-aware generative augmentation framework for virtual screening tasks. The SAS approach achieves particularly substantial improvements in EF100 and DCG100 metrics, with performance gains reaching approximately 45% and 38%, respectively, in optimal cases. These enrichment-based metrics are critical indicators of early retrieval performance in drug discovery workflows. The UA variant shows markedly smaller improvements for these same metrics and, more significantly, performs worse than the oversampling baseline for EF100 (declining over 10%) and DCG100 metrics when applied to certain model architectures. This performance degradation occurs because uniform sampling of active compounds worsens structural imbalance by generating additional molecules from already over-represented scaffold clusters. The resulting bias toward dominant structural patterns impairs the model’s ability to generalize to structurally distinct molecules during evaluation. These results underscore the importance of scaffold-aware sampling strategies that actively increase scaffold diversity among active compounds in the training data, rather than simply augmenting the total number of active samples.
Reranking Module
Note that we do not apply the reranking module for the results reported in Table 1. In Figure 3, we study the effect of the reranking module by varying the trade-off parameter for . corresponds to ScaffAug without reranking, where molecules are ordered only by the predicted scores, while activates the MMR module and gradually puts more weight on scaffold diversity. As decreases from , SD100 increases consistently, and EF100 stays similar or even improves in several assays, which shows that reranking adds clear value on top of the base model by improving scaffold diversity in the top hits without sacrificing early enrichment.